# Miura transformation in bidifferential calculus and a vectorial Darboux transformation for the Fokas-Lenells equation

### v1

@@ 1 Introduction
For the class of integrable equations possessing a “bidifferential calculus representation” [33], there is a universal vectorial Darboux transformation. Given an integrable equation from this class, it is usually straight forward to deduce the corresponding vectorial Darboux transformation from the universal one. Such an approach was taken in [34], where, however, the coupled Fokas-Lenells equations were only reached via a Miura transformation. [^3] In the present context, a “Miura transformation” is understood as a relation expressing the dependent variable of a differential equation in terms of the dependent variable of another differential equation, and its derivatives, implying that any solution of the latter is a solution of the former equation. It is named after Robert M. Miura, who discovered such a relation between the KdV and the modified KdV equation. As a consequence, the computations turned out to be overly complicated and the results obtained were not sufficiently general. By using observations in [35], more powerful results can be achieved in a much more elegant way, which also greatly improve those in [35], where the so-called Cauchy matrix approach has been taken to generate exact solutions.
@@ 3 Application to the coupled Fokas-Lenells equations
$\phi_{x}=\frac{1}{2}\,[J,\tilde{g}]\,\tilde{g}^{-1}\,,\qquad\tilde{g}_{t}\tilde{g}^{-1}+(\log\det(g))_{t}I_{2}=\frac{1}{2}\,[J,\phi]\,.$
Let now $J=\mathrm{diag}(1,-1)$. From the second equation, by taking the trace, we obtain
$(\log\det(g))_{t}=0\,,$
which reduces the Miura system (3.1) to
@@
Proposition 3.3.
Where $u$ and $v$ are non-zero, the Miura system (3.10) is equivalent to
@@
$\tilde{g}_{11}=\frac{1+\mathrm{i}\,u_{x}v}{\tilde{g}_{22}}\,,\qquad\tilde{g}_{21}=-\frac{u_{x}}{\tilde{g}_{22}}\,,$ (3.13)
$\phi_{12}=-\mathrm{i}\,v_{t}+uv^{2}\,,$ (3.14)
$\phi_{11x}=-\phi_{22x}=-\mathrm{i}\,u_{x}v\,,$ (3.15)
and the coupled Fokas-Lenells equations (1.2).
@@
$\tilde{g}_{22t}=\mathrm{i}\,u\,v\,\tilde{g}_{22}\,,\qquad\tilde{g}_{21}\,\tilde{g}_{22}=-u_{x}\,,\qquad u\,\tilde{g}_{11}=-\tilde{g}_{21t}\,,\qquad\tilde{g}_{12}=-\mathrm{i}\,v\,\tilde{g}_{22}\,,$
@@
The general solution of the first equation is given by (3.12). The last system then determines, in turn, $\tilde{g}_{21}$, $\tilde{g}_{11}$ (assuming $u\neq 0$) and $\phi_{12}$ in terms of $u$ and $v$. Inserting the results in the last equation of the above system, turns it into the first of (1.2). In the same way, the penultimate equation of the system becomes the second of (1.2), assuming $v\neq 0$. The remaining equations are in (3.15). ∎
It is easily verified that (3.12) - (3.15) is also a solution of (3.10) if $u=0$ or $v=0$. Hence, given any solution $(u,v)$ of the coupled Fokas-Lenells equations (1.2), (3.12) - (3.15) determines a solution $\tilde{g}$ of the Miura system (3.10). Next we present a vectorial binary Darboux transformation for the coupled Fokas-Lenells equations (1.2).
@@
with $n$-component row vectors $\theta_{i}$ and $n$-component column vectors $\eta_{i}$, $i=1,2$, and using the solution of the Miura system corresponding to $(u,v)$ according to Proposition 3.3, this becomes (3.16). By an application of (1.2),
$\phi_{12x}=-\mathrm{i}\,v\,(1+\mathrm{i}\,u_{x}v)\,.$

### v2

@@ 1 Introduction
For the class of integrable equations possessing a “bidifferential calculus representation” [33], there is a universal vectorial Darboux transformation. Given an integrable equation from this class, it is usually straight forward to deduce the corresponding vectorial Darboux transformation from the universal one. Such an approach was taken in [34], where, however, the coupled Fokas-Lenells equations were only reached via a Miura transformation. [^3] In the present context, a “Miura transformation” is understood as a relation expressing the dependent variable of a differential equation in terms of the dependent variable of another differential equation, and its derivatives, implying that any solution of the latter determines a solution of the former equation. It is named after Robert M. Miura, who discovered such a relation between the KdV and the modified KdV equation. As a consequence, the computations turned out to be overly complicated and the results obtained were not sufficiently general. By using observations in [35], more powerful results can be achieved in a much more elegant way, which also greatly improve those in [35], where the so-called Cauchy matrix approach has been taken to generate exact solutions.
@@ 3 Application to the coupled Fokas-Lenells equations
$\phi_{x}=\frac{1}{2}\,[J,\tilde{g}]\,\tilde{g}^{-1}\,,\qquad\tilde{g}_{t}\tilde{g}^{-1}+\frac{(\sqrt{\det(g)})_{t}}{\sqrt{\det(g)}}\,I_{2}=\frac{1}{2}\,[J,\phi]\,.$
From the second equation, by taking the trace, we obtain
$\det(g)_{t}=0\,.$
Let now $J=\mathrm{diag}(1,-1)$. Then the Miura system (3.1) results in
@@
Lemma 3.3.
(1) If $u$, $u_{x}$ and $v$ are non-zero, the Miura system (3.10) implies
@@
$\tilde{g}_{11}=\frac{1+\mathrm{i}\,u_{x}v}{\tilde{g}_{22}}\,,\qquad\tilde{g}_{12}=-\mathrm{i}\,v\,\tilde{g}_{22}\,,\qquad\tilde{g}_{21}=-\frac{u_{x}}{\tilde{g}_{22}}\,,$ (3.13)
$\phi_{12}=-\mathrm{i}\,v_{t}+uv^{2}\,,\qquad\phi_{21}=u\,,$ (3.14)
$\phi_{11x}=-\phi_{22x}=-\mathrm{i}\,u_{x}v\,.$ (3.15)
(2) If $(u,v)$ solves the coupled Fokas-Lenells equations (1.2), then (3.12)-(3.15) determines a solution $(\phi,\tilde{g})$ of the Miura system (3.10).
@@
$\tilde{g}_{22t}=\mathrm{i}\,u\,v\,\tilde{g}_{22}\,,\qquad\tilde{g}_{21}\,\tilde{g}_{22}=-u_{x}\,,\qquad u\,\tilde{g}_{11}=-\tilde{g}_{21t}\,,\qquad\phi_{12}=-\mathrm{i}\,v_{t}+u\,v^{2}\,,$
@@
The general solution of the first equation is given by (3.12) and everywhere non-zero on $\mathbb{R}^{2}$.
(1) Let (3.10) hold. According to Proposition 3.1, $u$ and $v$ then satisfy (1.2). The last system determines, in turn, $\tilde{g}_{12}$, $\tilde{g}_{21}$, $\tilde{g}_{11}$ (assuming $u\neq 0$ and using (1.2)) in terms of $u$ and $v$. The obtained expressions are in (3.13). The last two equations of the above system are then satisfied as a consequence of (1.2), and the remaining equations are given by (3.15).
(2) This is easily verified. ∎
Next we present a vectorial binary Darboux transformation for the coupled Fokas-Lenells equations (1.2).
@@
with $n$-component row vectors $\theta_{i}$ and $n$-component column vectors $\eta_{i}$, $i=1,2$, and using the solution of the Miura system corresponding to $(u,v)$ according to Lemma 3.3, this becomes (3.16). We note that $\phi_{12x}=-\mathrm{i}\,v\,(1+\mathrm{i}\,u_{x}v)$.